6<-4x+5 or 9<=-5x+3
What is the solution to the compound inequality.

Answer by lwsshak3(11628)

6<-4x+5 or 9<=-5x+3
What is the solution to the compound inequality
..
6<-4x+5
-4x+5>6 (I like to turn it around)
-4x>6-5=1
-x>1/4
x<-1/4
..
9<=-5x+3
-5x+3>9
-5x>9-3=6
-x>6/5
x<-6/5
..
Since this is an "and" function, the solution set is the intersection or overlap, that is, both inequalities must be true at the same time.
ans:
solution set(-∞,-6/5)